# Community Forensics: Using Thousands of Generators to Train Fake Image Detectors

### v1

@@ 4.1 Training image forensics models
Figure 3: Performance across generator types. We evaluate the classifier performance across five generator types – latent diffusion, pixel diffusion, commercial models, GANs, and other architecture (Stable Cascade [97]). Our classifiers show robust performance across all generator types, where prior works struggle to generalize.
4.2 Generalization to other datasets
We first evaluate how well classifiers trained on our dataset transfer to other benchmarks. In Table 2, we observe that our models outperform the prior works [91, 126, 133] in all evaluation sets except GenImage. This is expected since the GenImage evaluation set is a validation split of their training set; all of the generators are already seen by their classifier. On all unseen evaluation sets, our classifiers outperform all prior works. Notably, our classifiers achieve very high performance (0.991 mAP and 92.5% accuracy) on our comprehensive evaluation set with a significant margin compared to prior works. This gap in performance can be traced to our training data which incorporates a substantially richer variety of generators compared to prior works. Consequently, our classifiers demonstrate robust generalization to out-of-distribution data, where prior works often struggle.
To better illustrate the generalization of the classifiers, we show the performance per each generator type in Figure 3. We group them into five subsets: latent diffusion, pixel diffusion, commercial models, GANs, and other architecture type (Stable Cascade [97]). Our classifiers show strong performance across all generator types, unlike prior works which struggle to generalize to diverse architectures.
For the following experiments, we use our best-performing model (High res.) unless stated otherwise.
Figure 4: Performance with increasing number of images. We train a classifier with varying numbers of images from two sets: 1000 randomly chosen models and 10 popular (highly downloaded) models in the systematically collected subset. The classifier trained from 1000 random models outperforms 10 popular models in all cases. Notably, the accuracy gap is wider than that of mAP, which may suggest that having diversity in the models improves accuracy threshold calibration. We report the mean and standard error values for each data point across 4 randomly sampled subsets.
@@ 4.3 Impact of model diversity
Next, we examine the impact of the number of models in training data. We train classifiers with images sampled from 3 to $3$ generators and evaluate them (Fig. 1). To ensure that the gains are not due to simply sampling qualitatively different architectures, we only use our systematically collected latent diffusion models. We use an extended evaluation set that includes non-latent diffusion generators from our training set, which allows us to comprehensively assess the generalization capability of the classifiers trained exclusively on latent diffusion models. We find that the performance steadily increases with the number of models. However, the performance begins to flatten out beyond 1000 models, suggesting diminishing returns. Interestingly, the performance improves even on out-of-distribution architectures such as GANs and pixel-based diffusion models, even though the classifier is only trained on latent diffusion models.
@@
Our experiments show that the performance improvements from increasing the number of models may plateau when they are limited to a single generator type (Fig. 1). In Figure 5(a), we show that the diversity of the generator type also plays a major role in generalization. We train classifiers on three different sets of training data: our systematically collected set, manually chosen set, and a full set consisting of both subsets. The systematic set comprises entirely of latent diffusion models, and the manual set contains numerous generator types, including GANs, latent and pixel-based diffusion, and autoregressive models (Sec. 3.2). The classifier trained on the manual set with more diverse generator types shows stronger performance compared to the one trained on the systematic set. Additionally, we find that the two sets are complementary; the performance is further improved when we train using both sets.
@@ 4.4 Analysis of design choices
(a) Impact of frozen backbones
(b) Semantic alignment analysis
Figure 6: (a) Evaluating frozen backbones. Freezing the pretrained backbone, a common practice in prior works ([91, 18]), consistently decreases the performance. (b) Analyzing source and generated data alignment. We evaluate how the pairing of the real datasets affects performance. R denotes the real dataset used in training, and F indicates the source dataset used to obtain the captions for prompting the generators. The results suggest that pairing the source data (i.e., real data used to prompt the generators) with the generated images is not essential for performance.
@@
Prior works [91, 18] suggested using a frozen CLIP-ViT backbone for training the classifiers. We investigate this practice by training the classifiers with both frozen and unfrozen pretrained backbones, using CLIP-ViT [30, 100, 55] and ConvNeXt [78]. As shown in Figure 6(a), freezing the backbone consistently leads to poorer performance, indicating that end-to-end training is crucial to achieving high performance.
@@
Existing works often pair the generated images with the source dataset (i.e., the real dataset used to prompt or generate the images) arguing that misaligned data can introduce bias [126, 91, 18, 9]. We test this practice in Fig. 6(b) by examining the performance with both semantically aligned and misaligned real datasets. Specifically, we consider two real datasets: one comprised exclusively from LAION [110] and another combining ImageNet [26], MS-COCO [75], LandscapesHQ [115], Forchheim [45], VISION [114], and IMD2020 [90]. We sample our systematically collected latent diffusion models using these two sets and categorize the generated images by their source dataset. The resulting changes in the performance of all pairs were marginal, suggesting that strict alignment may not be as critical as previously believed.
@@
5 Discussion
In this paper, we studied the problem of generalizing to unseen generative models in synthetic image detection. We proposed a new dataset, Community Forensics, which contains 4803 models and 2.7M images collected from various public sources. We studied the impact of model diversity on the generalization performance of the classifiers and demonstrated that diverse data is a crucial factor in training a robust, generalizable forensics model. We trained classifiers on our dataset, demonstrated their ability to generalize on various settings, and evaluated previously proposed architecture and training practices.
While we only focus on generated image detection in our paper, our dataset may enable further forensics studies that can take advantage of our diverse array of generators. For example, in Appendix A, we show a preliminary result in identifying the generator used to synthesize a given image.
@@ 5 Discussion
While our dataset is diverse, a large portion of the data is diffusion-based, especially models fine-tuned on Stable Diffusion [102]. However, we note that Figure 1 shows that classifiers trained only on diffusion models still generalize reasonably well to other generator types as they represent different semantic content. Future work may consider collecting more diverse models, including more GANs, VQ-VAEs [123], and autoregressive models. We also note that the generative models sourced from the community may contain inappropriate content. While in many contexts it is important to detect such images (and we have removed images flagged by a NSFW detector), these models may require further scrutiny before being used in other downstream applications. Finally, although our experiments suggest that our forensics classifiers generalize to unseen models better than those of previous work, the error rates for forensics models are still too high to be used in many important applications.
@@ Acknowledgements.
We thank the creators of the many open source models that we used to collect the Community Forensics dataset. We thank Chenhao Zheng, Cameron Johnson, Matthias Kirchner, Daniel Geng, Ziyang Chen, Ayush Shrivastava, Yiming Dou, Chao Feng, Zihao Wei, Zixuan Pan, Inbum Park, Rohit Banerjee, and Ang Cao for the valuable discussions and feedback. This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001120C0123.
@@ Appendix B Dataset composition
Table 4: Model counts per architecture in the training set. A vast majority of the generators are latent diffusion models.
Figure 11: Number of images per generator type in the training set.
Figure 12: Evaluation set composition.
@@ Appendix D Example model project page
Figure 14: Example model project page from Hugging Face [2, 37].
Figure 14 shows a project page from Hugging Face [37, 2]. We can see the tags associated with the model (e.g., Text-to-image, pipeline type, license), number of downloads, and sample images.

### v2

@@ 4.1 Training image forensics models
Figure 3: Performance across generator types. We evaluate the classifier performance across five generator types – latent diffusion, pixel diffusion, commercial models, GANs, and other architecture (Stable Cascade [97]), and report the mean. We additionally evaluate three forensics methods, Aeroblade [103], LGrad [120], and NPR [121]. Our classifiers show robust performance across all generator types, whereas prior works struggle to generalize.
@@
Figure 4: Performance with increasing number of images. We train a classifier with varying numbers of images from two sets: 1000 randomly chosen models and 10 popular (highly downloaded) models in the systematically collected subset. We report the mean and standard error bands for each data point across 4 randomly sampled subsets. The classifier trained from 1000 random models outperforms 10 popular models in all cases. Notably, the accuracy gap is wider than that of mAP, which may suggest that having a diverse set of models improves accuracy threshold calibration.
4.2 Generalization across benchmarks
We first evaluate how well classifiers trained on our dataset perform across benchmarks. In Table 2, we observe that our models outperform the prior works [90, 129, 137, 43] on all evaluation sets except GenImage. This is expected since the GenImage evaluation set is a validation split of their training set; all of the generators are already seen by their classifier. On all other unseen benchmarks, our classifiers outperform all prior works. Notably, our classifiers achieve very high performance (0.987 mAP and 89.2% accuracy) on our out-of-distribution, comprehensive evaluation set, with a significant margin over prior works. This gap in performance can be traced to our training data which incorporates a substantially richer variety of generators than existing works. Consequently, our classifiers demonstrate robust generalization to out-of-distribution data, where prior works often struggle.
To better illustrate the generalization of the classifiers, in Figure 3, we split our comprehensive evaluation set into five subsets and evaluate them: latent diffusion, pixel diffusion, commercial models, GANs, and other architecture type (Stable Cascade [97]). Furthermore, we evaluate three additional forensics methods: Aeroblade [103], LGrad [120], and NPR [121]. Our classifiers show strong performance across all generator types, unlike prior works which struggle to generalize. For the following experiments, we use our best-performing model (High res.) unless stated otherwise.
@@ 4.3 Impact of model diversity
Next, we examine the impact of the number of models in training data. We train classifiers with images sampled from 3 to $3$ generators and evaluate them (Fig. 1). To ensure that the gains are not due to simply sampling qualitatively different architectures, we only use our systematically collected latent diffusion models. For each data point, we sample 10 random subsets of models with 100K training images each and report the mean and standard error bands. We use an extended evaluation set that includes non-latent diffusion generators from our training set, which allows us to comprehensively assess the generalization capability of the classifiers trained exclusively on latent diffusion models. We find that the performance steadily increases with the number of models. However, the performance begins to flatten out beyond 1000 models, suggesting diminishing returns. Interestingly, the performance also improves on out-of-distribution architectures such as GANs and pixel-based diffusion models, even though the classifier is only trained on latent diffusion models.
@@
Our experiments show that the performance improvements from increasing the number of models may plateau when they are limited to a single generator type (Fig. 1). In Figure 5(a), we show that the diversity of the generator type also plays a major role in generalization. We train classifiers on three different sets of training data: our systematically collected set, manually chosen set, and a full set consisting of both subsets. The systematic set comprises entirely of latent diffusion models, and the manual set contains numerous generator types, including GANs, latent and pixel-based diffusion, and autoregressive models (Sec. 3.2). The classifier trained on the manual set with more diverse generator types shows similar performance compared to the one trained on the systematic set, despite containing fewer than half the number of images (774K vs. 1.9M). Additionally, we find that the two sets are complementary; the performance is further improved when we train using both sets.
(a) Impact of frozen backbones
(b) Semantic alignment analysis
Figure 6: (a) Evaluating frozen backbones. Freezing the pretrained backbone, a common practice in prior works [90, 15], consistently decreases the performance. (b) Analyzing source and generated data alignment. We evaluate how the pairing of the real datasets affects performance. ‘R’ denotes the real dataset used in training, and ‘F’ indicates the source dataset used to obtain the captions for prompting the generators. The results suggest that pairing the source data (i.e., real data used to prompt the generators) with the generated images is not essential for performance.
@@ 4.4 Analysis of design choices
Prior works [90, 15] suggested using a frozen CLIP-ViT backbone for training the classifiers. We study this by training the classifiers with both frozen and unfrozen pretrained backbones, using CLIP-ViT [27, 100, 54] and ConvNeXt [77]. As shown in Fig. 6(a), freezing the backbone consistently leads to worse results, indicating that end-to-end training is crucial for high performance.
@@
Existing works often pair the generated images with the source dataset (i.e., the real dataset used to prompt or generate the images) arguing that misaligned data can introduce bias [129, 90, 15, 7]. We test this practice in Fig. 6(b) by examining the performance with both semantically aligned and misaligned real datasets. Specifically, we consider two real datasets: one comprised exclusively from LAION [111] and another combining ImageNet [23], MS-COCO [74], LandscapesHQ [116], Forchheim [44], VISION [115], and IMD2020 [89]. We sample our systematically collected latent diffusion models using these two sets and categorize the generated images by their source dataset. The performance differences of all pairs were marginal, suggesting that strict alignment may not be as critical with sufficiently large data.
@@
4.5 Other forensics applications
Our dataset may enable further forensics studies that can take advantage of our diverse array of generators. To illustrate this, in Appendix A, we identify the type of generator used to synthesize a given image by using $k$-nearest neighbor in the feature space of our classifier.
5 Conclusion
In this paper, we studied the problem of generalizing to unseen generative models in synthetic image detection. We proposed a new dataset, Community Forensics, which contains 4803 models and 2.7M images collected from various public sources. We studied the impact of model diversity and demonstrated that it plays a crucial role in enhancing data diversity and generalization performance. We trained classifiers on our dataset, studied their ability to generalize in various settings, and evaluated previously proposed models and training practices.
@@ 5 Conclusion
While our dataset is diverse, a large portion of the data is diffusion-based, especially models based on Stable Diffusion [104]. Despite this, Figure 1 demonstrates that adding more diffusion models to the training data improves the generalization capability of the classifiers across other generator types. Furthermore, Figure 5(a) illustrates improved classifier performance when combining the diffusion-only systematic set with a more diverse manual set. These results suggest that despite their similarities, each diffusion model may uniquely contribute through its semantic content, image processing pipelines, or architectural variations (Sec. 3.1). Future work may explore collecting more diverse models, including GANs, VQ-VAEs [126], and autoregressive models. We also note that the generative models sourced from the community may contain inappropriate content. While in many contexts it is important to detect such images, these generated images may require further scrutiny before being used in other downstream applications. Finally, although our experiments suggest that our forensics classifiers generalize to unseen models better than those of previous work, their error rates are still too high for many critical applications.
@@ Acknowledgements.
We thank the creators of the many open source models that we used to collect the Community Forensics dataset. We thank Chenhao Zheng, Cameron Johnson, Matthias Kirchner, Daniel Geng, Ziyang Chen, Ayush Shrivastava, Yiming Dou, Chao Feng, Xuanchen Lu, Zihao Wei, Zixuan Pan, Inbum Park, Rohit Banerjee, and Ang Cao for the valuable discussions and feedback. This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001120C0123.
@@ Appendix B Dataset composition
Table 4: Model counts per architecture in the training set. The generators are predominantly latent diffusion models.
Figure 11: Number of images per generator type in the training set.
@@ Appendix D Example model project page
Figure 14: Example model project page from Hugging Face [34].
Figure 14 shows a project page from Hugging Face [34]. We can see the tags associated with the model (e.g., Text-to-image, pipeline type, license), number of downloads, and sample images.